# SATURN: SAT-based Reinforcement Learning to Unleash Language Model Reasoning

### v1

@@ 6 Conclusion and Future Work
  • Appendix C: Derivation of the SAT Difficulty Estimation
@@
  • Appendix I: Multi-stage SATURN Results on Math and Programming Tasks
  • Appendix J: Behavior of Stronger LLMs on SATURN Tasks
  • Appendix K: Examples of LLMs Reasoning Trajectories
  • Appendix L: Comparisons of Existing Reasoning Tasks
  • Appendix M: SATURN-7B Word Cloud on GPQA Diamond
@@ Appendix A Pseudocode of SATURN Algorithm and Hyperparameters
Algorithm 1 SATURN Learning_Loop$(n,k,l,\pi_{\theta})$  # LLM represents $\pi_{\theta}$
@@ Appendix B Construction Procedure of SAT Instances
[…]
@@
[…]
@@
Appendix C Derivation of Difficulty Estimation for SAT Tasks on LLMs
@@
where we approximate $12^{n}-$ 1 by $2^{n}$ for analytical simplicity.
@@
Second, the clause length $l$ determines the number of symbols each clause contains. Longer clauses introduce more intra-clause dependencies, increasing local reasoning complexity. We approximate this with:
@@
These two components affect reasoning difficulty independently—one globally through variable sharing, the other locally through clause complexity. We therefore combine them additively into:
@@
This appendix provides additional details on the SATURN-1.5B and SATURN-7B training. We conduct all experiments on 8 NVIDIA A100 40GB GPUs. We use the OpenRLHF framework [^1] https://github.com/OpenRLHF/OpenRLHF for GRPO training. The framework is designed to make RL training simple and user-friendly, and works well in our experiments. The hyperparameters used in training are summarized in Table 6. All other parameters not listed above consistently follow the default settings of OpenRLHF.
@@
Initial KL coefficient $1\times 10^{-3}$ $1\times 10^{-3}$
@@ Appendix E Evaluation Hyperparameters
Appendix F Evaluation Prompt Templates
This appendix provides the prompt templates used for evaluation, ensuring consistency and reproducibility across tasks. Figure 8 presents the format for SAT problem evaluation, while Figure 9 shows the template used for math, programming, and GPQA Diamond tasks.
Figure 8: Prompt format used for SAT problem evaluation.
@@ Appendix H Ablation Studies for SATURN
[…]
Table 13: Performance Comparison on Math and Programming Benchmarks
@@
Appendix I Multi-stage SATURN Results on Math and Programming Tasks
This appendix presents detailed results of SATURN-1.5B and 7B on math and programming benchmarks, with the performance summarized in Table 15. As the number of SATURN curriculum learning stages increases, the performance of LLMs improves, but the extent of improvement gradually diminishes and eventually reaches saturation.
Model AIME 24/25 AMC 22/23 Math500 GPQA-D LiveCodeBench Avg.
DeepSeek-R1-Distill-Qwen-1.5B 21.6 65.1 83.6 30.3 16.4 43.4
SATURN-1.5B-Iteration-1 26.7 68.6 85.0 33.3 16.9 46.1
SATURN-1.5B-Iteration-2 28.3 73.5 84.6 37.4 17.4 48.2
SATURN-1.5B-Iteration-3 28.3 66.3 85.8 36.9 16.7 46.9
DeepSeek-R1-Distill-7B 50.0 80.7 93.2 49.0 35.4 61.7
SATURN-7B-Iteration-1 48.3 83.1 94.6 50.5 36.6 62.6
SATURN-7B-Iteration-2 48.3 85.5 95.0 50.5 37.7 63.4
SATURN-7B-Iteration-3 46.7 87.9 93.2 58.1 38.1 64.8
Appendix J Behavior of Stronger LLMs on Extended SATURN Tasks
@@ Appendix I Multi-stage SATURN Results on Math and Programming Tasks
Appendix K Examples of Different LLMs’ Reasoning Trajectories
@@
Appendix L Detailed Comparison of Constructed Reasoning Tasks
Figure 14: Comparison of Knights and Knaves (K&K) logic puzzle and Boolean Satisfiability (SAT) Problem.
@@
K&K puzzles require translating symbolic logic into natural language. This involves paraphrasing logical constraints into grammatically and semantically valid sentences, which increases construction cost. In contrast, SAT problems are purely symbolic and follow a standard format. As shown in [18], symbolic SAT and its natural language version yield similar reasoning performance for LLMs, suggesting the symbolic form is sufficient.
@@
Appendix M Word Cloud of SATURN-7B’s Outputs on GPQA Diamond

### v3

@@ 80. Guidelines:
  • Appendix C: Derivation and Validity of the SAT Difficulty Estimation
@@
  • Appendix I: Behavior of Stronger LLMs on SATURN Tasks
  • Appendix J: Examples of LLMs Reasoning Trajectories
  • Appendix K: Comparisons of Existing Reasoning Tasks
  • Appendix L: SATURN-7B Word Cloud on GPQA Diamond
@@ Appendix A Pseudocode of SATURN Algorithm and Hyperparameters
[…]
@@ Appendix B Construction Procedure of SAT Instances
    Under the constraint of satisfying the solution, the $n$ variables are randomly negated, resulting in $2^{n}-1$ satisfiable clauses per variable set. From these steps, we uniformly obtain a large number of single-clause samples from the total of $2^{k}\cdot\binom{k}{n}\cdot(2^{n}-1)$ valid $(\texttt{solution},\texttt{clause})$ pairs, with an upper bound set to 100,000100{,}000.
@@
    These $(\texttt{solution},\texttt{clause})$ pairs are grouped into clusters based on the same solution.
@@
Appendix C Derivation and Validity of Difficulty Estimation for SAT Tasks on LLMs
@@
where we approximate $2^{n}-1$ by $2^{n}$ for analytical simplicity.
@@
Second, the clause length $l$ determines the number of symbols that each clause contains. Longer clauses introduce more intra-clause dependencies, increasing local reasoning complexity. We approximate this with:
@@
These two components affect reasoning difficulty independently—one globally through variable sharing, and the other locally through clause complexity. We, therefore, combine them additively into:
@@
Step 4: Ablation Study about estimation metric
To further validate the necessity and effectiveness of Eq. (12) composite metric, we conduct an ablation study comparing alternative formulations. Table 7 reports the $R^{2}$ correlation of each metric with LLM performance (pass@3) across multiple model scales.
Table 7: Metric comparison across models.
Metric Formula R1-Qwen-1.5B Saturn-1.5B R1-Qwen-7B Saturn-7B Avg. Std. Dev.
$k-l\cdot\log_{2}(1-\frac{1}{2^{n}})$ 0.507 0.537 0.132 0.000 0.294 0.269
$\alpha\cdot\frac{k}{n}+\beta\cdot\log_{2}(l),\ \alpha=2,\ \beta=1$ 0.428 0.478 0.568 0.508 0.496 0.059
$\alpha\cdot\frac{k}{n}+\beta\cdot\log_{2}(l),\ \alpha=1,\ \beta=1$ 0.240 0.279 0.719 0.826 0.516 0.300
$\log_{2}(k\cdot l)-\log_{2}(2^{n}-1)$ 0.875 0.893 0.451 0.157 0.594 0.356
$\log_{2}(k)+2\cdot\log_{2}(l)-n+\frac{k}{n}$ 0.707 0.746 0.724 0.501 0.670 0.113
Simpler metrics that consider only sparsity or only structural complexity perform worse overall, confirming that both components are essential for accurately capturing task difficulty. Our metric $\log_{2}(k)+2\cdot\log_{2}(l)-n+\frac{k}{n}$ combines both solution sparsity and structural complexity, which achieves the best overall correlation. The ratio between the solution space and the LLM’s search space is the most crucial aspect. On SATURN‑7B, the $R^{2}$ value is about 0.5 because the LLM already achieves over 90% accuracy on easy problems, which limits the observable linear correlation in that range. The corresponding figures are shown in Figures 10 and 11.
@@
This appendix provides additional details on the SATURN-1.5B and SATURN-7B training. We conduct all experiments on 8 NVIDIA A100 40GB GPUs. We use the OpenRLHF framework [^1] https://github.com/OpenRLHF/OpenRLHF for GRPO training. The framework is designed to make RL training simple and user-friendly, which works well in our experiments. The hyperparameters used in training are summarized in Table 8. All other parameters not listed above consistently follow the default settings of OpenRLHF.
@@
Initial KL coefficient $\times 10^{-3} 1×1$$\times 10^{-3}$
@@ Appendix E Evaluation Hyperparameters
Appendix F Prompt Templates
This appendix provides the prompt templates used for evaluation, ensuring consistency and reproducibility across tasks. Figure 8 presents the format for SAT problem training and evaluation, while Figure 9 shows the template used for math, programming, and GPQA Diamond tasks.
Figure 8: Prompt format used for SAT problem training and evaluation.
@@ Appendix H Ablation Studies for SATURN
This appendix presents the ablation studies for SATURN, as shown in Table 15 and Table 16. Each training setting is denoted as $(n,k,l)\times\texttt{Train\_size}$, where $(n,k,l)$ controls SAT construction and Train_size is the number of training examples. Here, Train_size can also be written as Train_size = $D\times\texttt{num}$, where $D$ is the number of difficulty levels and num is the number of samples per level in SATURN-2.6k (Figure 11 and 10). These experiments validate the effectiveness of curriculum learning and the design of various SAT training configurations.
Table 15: Ablation comparison on math and programming benchmarks
@@
Appendix I Behavior of Stronger LLMs on Extended SATURN Tasks
@@
To provide a baseline comparison, we also tested a CDCL SAT solver  [^5] https://en.wikipedia.org/wiki/Conflict-driven_clause_learning on the SATURN-2.6k test set. The results are as follows:
Table 18: CDCL SAT solver performance on SATURN-2.6k test set.
Metric Value
Total instances 1000
Satisfiable 1000
Unsatisfiable 0
Valid SAT Models 1000
Model Accuracy 100.00%
Total time taken (s) 0.14
Table 18 shows that the CDCL SAT solver significantly outperforms current LLMs like DeepSeek-R1, both in terms of runtime and accuracy. SATURN applied to today’s strongest LLMs still has great potential.
Appendix J Examples of Different LLMs’ Reasoning Trajectories
@@
Appendix K Detailed Comparison of Constructed Reasoning Tasks
Figure 14: Comparison of Knights and Knaves (K&K) logic puzzle and SAT problem.
@@
K&K puzzles require translating symbolic logic into natural language. This involves paraphrasing logical constraints into grammatically and semantically valid sentences, which increases construction cost. In contrast, SAT problems are purely symbolic and follow a standard format. As shown in [sat_iclr_1], symbolic SAT and its natural language version yield similar reasoning performance for LLMs, suggesting the symbolic form is sufficient.
@@
Appendix L Word Cloud of SATURN-7B’s Outputs on GPQA Diamond